OpenStudy (anonymous):

Solve for x: x + 16 / 3 = 3x

OpenStudy (anonymous):

What Do You Think Is First? Divide?

OpenStudy (mayankdevnani):

-2x=-16/3 x=16/3*2

OpenStudy (mayankdevnani):

can you solve it

OpenStudy (anonymous):

x + 16 / 3 = 3x -2x = 16/3 x = -16/6

OpenStudy (anonymous):

x + 16 / 3 = 3x -x -x 16/3=2x /2 /2 8/3=x

OpenStudy (mayankdevnani):

@gerryliyana warning! don't give direct answers

OpenStudy (farmdawgnation):

@ gerryliyana You should heed the advice of your peers. Do just give out "just answers". Thanks.

OpenStudy (lgbasallote):

do?

OpenStudy (farmdawgnation):

Erhm. * Don't

OpenStudy (farmdawgnation):

Haven't had my coffee yet.

OpenStudy (anonymous):

she needs to study timetable =|

OpenStudy (anonymous):

Im horrible in math....

OpenStudy (farmdawgnation):

@Stephhaniiee Then allow the people here to help you get better. That's what we're here for. We don't want to do your homework for you. Yahoo! Answers can do that. We want to teach you. Let the people on this site do that.

OpenStudy (anonymous):

i am sorry ...,

OpenStudy (anonymous):

so who agrees that the answer is 8/3=x

OpenStudy (anonymous):

i am

OpenStudy (anonymous):

i am agreed

OpenStudy (farmdawgnation):

Is the problem \(x + \frac{16}{3}\) or \(\frac{x + 16}{3}\)?

OpenStudy (farmdawgnation):

Sorry, left off the RHS there. heh.

OpenStudy (farmdawgnation):

Oof. I need to finish my abuse report audit. Sorry guys. You're on your own. :)

OpenStudy (farmdawgnation):

@lgbasallote knows all the things.